Determine whether the statement below is true or false. If it is false, rewrite it as a true statement Choose the correct answer below. O A. This statement is true. O B. This statement is false. A true statement is 7C4-7C2. O C. This statement is false. A true statement is C5 7P2 0 D. This statement is false. A true statement is 2C7-5C7.
Answer #1

Solution :

We have to check it .

nCx = n! / x!(n - x)!

n! = n(n - 1)(n - 2) .... and so on .

n = 7 and x = 5

7C5 = 7! / 5!(7 - 5)! = 7! / 5!(2)! = 7 * 6 / 2 * 1 = 42 / 2 = 21

n = 7 and x = 2

7C2 = 7! / 2!(7 - 2)! = 7! / 2!(5)! = 7 * 6 / 2 * 1 = 42 / 2 = 21

A) This statement is true .